A suggested mechanism for the decomposition of ozone is as follows: O3 O2 + O fast equilibrium O + O3 2O2 slow step What is the rate law predicted by this mechanism? 


1
Expert's answer
2021-09-26T12:14:58-0400

For a reaction, mechanism follows the next elementary steps:



Here, intermediate is O as it forms in one elementary step and it is consumed in next steps.

Rate law comes from slowest step that is,

"rate = k_2[O][O_3] \;\;\;(1)"

For fast step at equilibrium, rate of forward and reverse reactions are equal.

"k_1[O_3] =k_{-1}[O_2][O] \\

[O] = \frac{k_1[O_3]}{k_{-1}[O_2]} \;\;\;(2)"

Substitute [O] in rate law (1).

"rate = k_2[O][O_3] \\

= k_2\frac{k_1[O_3]}{k_{-1}[O_2]}[O_3] \\

= \frac{k_1k_2[O_3]^2}{k_{-1}[O_2]}"